WRITE AN EQUATION OF A LINE CONTAINING THE GIVEN POINT, AND PERPENDICULAR TO A GIVEN LINE: write answer in y=mx+b form
(9,2)
5x+y=7

Answer by nerdybill(7384)

Begin by finding the slope of:
5x+y=7
do this by putting the equation in "slope-intercept" form:
y = -5x + 7
by inspection, we see that the slope is -5
.
A line perpendicular, must have a slope that is the "negative reciprocal":
(-5)m = -1
m = 1/5 (this is the slope of the new line)
using the slope and the given point:
(9,2)
plug into the "point-slope" form:
y - y1 = m(x - x1)
y - 2 = (1/5)(x - 9)
y - 2 = (1/5)x - 9/5
y = (1/5)x - 9/5 + 2
y = (1/5)x - 9/5 + 10/5
y = (1/5)x + 1/5 (this is what they're looking for)